OpenStudy (mcba1234):

Mason and Nora decided to swim across the river. Mason began swimming 35 seconds earlier than Nora. Mason swam at a speed of 5 feet per second. Nora swam at a speed of 6 feet per second. For how many seconds had Mason been swimming at the moment when the two swimmers had swam exactly the same distance?

OpenStudy (welshfella):

speed = distance / time

OpenStudy (welshfella):

let d be distance travelled when they are level , and t = time that nora tokk to get level with mason for Nora 6 = d/t d = 6t for Mason 5 = d / (t + 35) d = 5 ( t + 350 so 6t = 5(t + 35) solve for t

OpenStudy (welshfella):

Masons time in the water will be t + 35